# Cascading from the winds to the disk: the universality of supernovae-driven turbulence in different galactic interstellar mediums

### v1

@@ 4.1 Definitions
$f^{2}(\bm{k})=|\tilde{\bm{f}}(\bm{k})\tilde{\bm{f}}^{\dagger}(\bm{k})|,$ (22)
where $\square^{\dagger}$ denotes the complex conjugate. Even though there is a global anisotropy imparted on the flow due to $\bm{\nabla}\phi$, it is never strong enough to make the $\ell\parallel\bm{\nabla}\phi$ and $\ell\perp\bm{\nabla}\phi$ statistics significantly different (Beattie et al., 2025c). Hence, in this study we integrate the three-dimensional spectrum isotropically over solid-angle shells where $\operatorname{d}\!{\Omega_{k}}$, where $k=|\bm{k}|$,
$f^{2}(k)=\int\operatorname{d}\!{\Omega_{k}}\;4\pi k^{2}f^{2}(\bm{k}),$ (23)
resulting in the one-dimensional isotropic spectrum, $f^{2}(k)$.
@@ 4.2 The $k$-space structure of the turbulence
In Figure 7, we plot the integrated spectra of the total velocity $\bm{u}$ (top row), the kinetic energy $\sqrt{\rho}\bm{u}$ (second row), the decomposed compressible $\bm{u}_{c}$ and incompressible $\bm{u}_{s}$ velocity modes (third row), and the sound speed $c_{s}$ (bottom row), for MW (left column) and SB (right column), all normalized by the integral power. The gaseous scale height, $k_{0}\sim\ell_{0}^{-1}$, is indicated with the grey vertical line.
@@
Beginning with the first row, we compensate the velocity spectra for Kolmogorov (1941)-type turbulence, and we find that all models have a shallower than the Kolmogorov (1941) spectrum, following the $k^{-3/2}$ spectrum found in the magnetohydrodynamic (MHD) SNe-driven turbulence study of Gent et al. (2021) and also the MHD studies by Beattie et al. (2025a) and Iroshnikov (1964). This spectrum has also been derived for weak acoustic wave turbulence (Zakharov & Sagdeev, 1970; Kochurin & Kuznetsov, 2024), and shown to potentially be embedded within supersonic turbulence if one completely resolves a subsonic cascade (Ferrand et al., 2020; Beattie et al., 2025a). This is plausible for our set up if SNe shocks become linear sound waves on small enough scales.
@@
Once we decompose each simulation into $\bm{u}_{c}$ and $\bm{u}_{s}$ (second row), we observe that, as found in a number of SNe-driven ISM simulations, e.g., Balsara et al. (2004), Padoan et al. (2016), Pan et al. (2016) and Beattie et al. (2025c), the turbulence is strongly dominated by the incompressible $\bm{u}_{s}$ modes, with a spectra shallower than $k^{-5/3}$, and especially so for the high-SN rate SB model. This aligns also with work that found incompressible mode dominance to be true even for purely compressible-driven simulations (Federrath, 2013), but likely the mechanism for generating the incompressible modes varies significantly between the simulations, as discussed in Beattie et al. (2025c), i.e., the shallower spectrum dominates at smaller scales than the scale height, even more so than in purely incompressible-driven turbulent boxes (Federrath, 2013). Beattie et al. (2025c) proposed that phase mixing through the fractal cooling layer of SNRs is the strongest source of incompressible turbulence in SNe-driven turbulence.
@@
The compressible mode spectrum in the second row closely follows Burgers (1948) $k^{-2}$ prediction for strongly nonlinear compressible turbulence for the SC seeding models, potentially with some shallowing towards $k^{-5/3}$ in the SBFX model, but generally, $k^{-2}$ without a break scale. We also plot the pseudo-acoustic wave spectrum $k^{-3}$, which has been realized in low-$\mathcal{M}$ ($\mathcal{M}\lesssim 0.2$) hydrodynamical turbulence (Wang et al., 2017; Beattie et al., 2025b), but as we see here, none of the simulations are in this regime.
@@
Because the turbulence is dominated by $\bm{u}_{s}$ modes it is unlikely that the $k^{-3/2}$ spectrum comes from the weak acoustic regime, and in order to get shallower spectra analogous to MHD turbulence, like $k^{-3/2}$ there must be a reduction in the turbulent nonlinearity (Boldyrev, 2006; Matthaeus et al., 2008; Chernoglazov et al., 2021; Beattie & Bhattacharjee, 2025), which in our case can possibly be attributed to the strong alignment between velocity $\bm{u}$ and vorticity $\bm{\omega}$ (the kinetic helicity), as seen in Beattie et al. (2025c) (in their Appendix, Figure 21) and previously in (Käpylä et al., 2018). The presence of an inverse cascade can also cause a shallower spectrum, as shown in detail by Beattie et al. (2025c) which is shown to be possible in helical, hydrodynamical turbulence (Plunian et al., 2020), potentially the kind of turbulence that is generated on the expanding SNRs.
@@
In the third row we plot the kinetic energy spectra, $\rho u^{2}(k)$, compensated by $k^{-4/3}$, found previously in compressible (magnetized) turbulence (Grete et al., 2021). For $k\lesssim\ell_{0}^{-1}$, we see close alignment between the $k^{-4/3}$ model and $\rho u^{2}(k)$. For SB, we see this behavior at small and large scales, but with a transition between at $k\approx\ell_{0}^{-1}$. This transition is most likely due to the gaseous profile of the disk imprinting itself across the spectrum. As discussed in the Appendix of Beattie et al. (2025c), this increase in kinetic energy density at high $k$, compared to the $u^{2}(k)$, is likely due to high-density clumps on small scales. The velocity power spectrum is more universal across scales and between models, suggesting the velocity modes are correlated through the turbulence in the winds and disk, which we will explore in more detail in Section 4.4.
@@ 4.3 What sets the large-scale phase structure in the ISM?
which means that, in the adiabatic limit where $t_{\rm turb}\ll t_{\rm cool}$, $c_{s}^{2}(k)\sim u_{c}^{2}(k)$, as we show in the bottom row [^2] In Appendix B we also show the relation between $u_{c}^{2}(k)$ and $(\bm{\nabla}\cdot\bm{u})^{2}(k)$, i.e., $u_{c}^{2}(k)\sim(\bm{\nabla}\cdot\bm{u})^{2}(k)/k^{2}\sim c_{s}^{2}(k)\propto k^{-2}$. of Figure 7. It should now make sense that $c_{s}^{2}(k)$ for the MW model follows almost exactly the $\bm{u}_{c}$ spectrum, given that the disk has a strong adiabatic component (Figure 6. For the wind-dominated SB model, at the disk scale and onward ($\ell_{0}$ annotated in black dashed line), the spectrum largely follows $k^{-2}$, since the SB disk is also largely adiabatic. For $k\lesssim\ell^{-1}_{0}$ the spectrum deviates, because the winds are not principally adiabatic and have a mix of thermodynamic processes that we do not account for in our simple adiabatic model. This shows the the important role of compressible turbulent modes in controlling the temperature and phases of the ISMs in our models.
Figure 8: One-dimensional, isotropic velocity (top) and kinetic energy (bottom) power spectrum, normalized into global gaseous scale heights, $k\ell_{0}/2\pi=1$, and power on the scale height (e.g., $u^{2}(k)/u^{2}(k\ell_{0}/2\pi)=1$) for each galaxy model (MW; dark purple, SB; light purple). We show different SNe seeding schemes (SC; dashed linestyle, FX; solid linestyle). The data that is greyed is either MW, where $k<\ell_{0}^{-1}$, or SB $k>\ell_{0}^{-1}$. By normalizing the spectrum into the frame of the scale height we find there is a smooth transition between both the velocity and kinetic energy for the turbulence in the winds and the turbulence in the disk. The velocity spectrum exhibits a universal power law, $u^{2}(k)\propto k^{-3/2}$, regardless of galaxy model or seeding prescription, whereas the kinetic energy spectrum retains some structure of the galactic disk, varying from $k^{-1/2}$ on $k<\ell_{0}^{-1}$ to $k^{-4/3}$ on $k>\ell_{0}^{-1}$. We indicate the global outer scale $\ell_{\rm{cor}}$ of the $u^{2}(k)$ and $\rho u^{2}(k)$ spectra (Equation 28) in the purple shaded line, showing (particularly for the $u^{2}(k)$ spectrum) that even though we inject momentum from SNe on the smallest scales in the simulations ($\ell_{\rm inj}=2\rm{dx}$), they drive turbulence to scales well beyond $\ell_{0}$, with $\ell_{\rm{cor}}\approx 6\ell_{0}$.
@@ 4.4 Cascading from the winds to the disk
Based on Figure 7, it is clear that many of the spectra seem to follow similar power law scalings across the different seeding and galactic models (which differ in mass, potential depth and injection energy). To explore this universality further, we transform the velocity and kinetic energy spectra such that $k\ell_{0}/2\pi=1$ for all models, and $u^{2}(k)/u^{2}(k\ell_{0}/2\pi)=1$, such that all models are placed in a global gaseous scale height reference frame. We plot the transformed spectra in Figure 8, showing the SB simulations in light purple (mostly situated in $k\lesssim\ell^{-1}_{0}$ modes, corresponding to the wind scales), and MW simulations in dark purple (mostly on the $k\gtrsim\ell^{-1}_{0}$ scales), and the two different seeding schemes with the different linestyles. We change the color of the spectra for the MW models on $k<\ell_{0}^{-1}$ using grey lines to indicate the spectra beyond those modes, and similarly for the SB spectra but for $k>\ell_{0}^{-1}$. This allows us to directly see how continuous the spectra are across $k\ell_{0}/2\pi$, whilst highlighting the structure of the turbulence in the different simulations below and above $k\ell_{0}/2\pi$.
We find that all spectra collapse across galactic models and seeding schemes into just a single velocity (top) and kinetic (bottom) spectrum, that are smooth across $\ell^{-1}_{0}$. The $u^{2}(k)\propto k^{-3/2}$ discussed in Section 4.2 and previously measured in Padoan et al. (2016) and Beattie et al. (2025c) extends across both simulations, connecting the winds to the turbulent disk in a continuous manner. For the $\rho u^{2}(k)$ spectrum we find a broken power law, that seems to break at roughly $k\ell_{0}/2\pi\approx 1$, with $\rho u^{2}(k)\propto k^{-1/2}$ on $k\ell_{0}/2\pi\lesssim 1$ and $\rho u^{2}(k)\propto k^{-4/3}$ on $k\ell_{0}/2\pi\gtrsim 1$. This is likely due to the gaseous disk imprinting itself on the power spectrum, due to the mass density contribution in the spectrum. Federrath (2013) previously found that the $\rho u^{2}(k)$ spectrum was sensitive to the underlying driving field and $\mathcal{M}$, whereas the $u^{2}(k)$ spectrum was not.
@@
$\frac{\ell_{\mathrm{cor}}}{\ell_{0}}=\frac{\displaystyle\int_{0}^{\infty}\operatorname{d}\!{k}\;(k\ell_{0}/2\pi)^{-1}f^{2}(k)}{\displaystyle\int_{0}^{\infty}\operatorname{d}\!{k}\;f^{2}(k)},$ (28)
in the global $\ell_{0}^{-1}$ frame using definitions from Beattie et al. (2025c), and find $\ell_{\rm cor}\approx 6\ell_{0}$ for $u^{2}(k)$, showing that the largest scale velocity correlations extend well beyond the gaseous scale-height, and $\ell_{\rm cor}\approx 3\ell_{0}$ for $\rho u^{2}(k)$, being truncated at small scales due to the imprint of the galactic disk on the $\rho u^{2}(k)$ spectrum. Both the spectra (but particularly the velocity spectrum), show that SNe can drive turbulence to large scales, larger than the conventional wisdom, $\ell_{0}\approx\ell_{\rm cor}$ (Haverkorn et al., 2008; Liu et al., 2022; Beattie et al., 2022a).
@@ 5 Summary and Conclusions
  • For each of the models we calculate the power spectrum of the total velocity, kinetic energy, Helmholtz decomposed velocity modes, and sound speed and plot them Figure 7. In the disk-dominated MW simulations we show the entire phase structure of the predominantly adiabatic disk can be explained by the $\bm{u}_{c}$ mode turbulence ($c_{s}^{2}(k)\propto u_{c}^{2}(k)\propto k^{-2}$), which we can derive by relating $c_{s}$ and $u_{c}$ in the weak-cooling, adiabatic limit, assuming that the compressible modes are sourced from turbulence within sound waves, detailed in Equations 24-27. Consistent with Beattie et al. (2025c), we find a velocity spectrum that is dominated by incompressible modes and close to $u^{2}(k)\propto k^{-3/2}$, suggesting a weakening of the turbulent nonlinearity which is robust to the different simulation parameters.
  • By transforming both the velocity $u$ and kinetic energy $\rho u^{2}(k)$ spectrum into units of $\ell_{0}$ and power on $\ell_{0}$, in Figure 8 we show that the wind-dominated (starburst; SB) and disk-dominated (Milky Way; MW) galaxy models have power spectra that smoothly transition into one another, suggesting the structure of the turbulence is robust to changes in the galactic potential, SNe-driving rate, and seeding prescription. Furthermore, a universal $u^{2}(k)\propto k^{-3/2}$ spectrum extends between the disks and the winds seamlessly, suggesting a wind-disk turbulence connection, with the velocity cascade moving through the different ISM phases without hindrance. Furthermore, in this frame, we show that the outer scale of the velocity is $\approx 6\ell_{0}$, challenging the conventional wisdom that the outer scale is truncated at the gaseous scale height of the disk. The kinetic energy spectrum traces large-scale structure from the galactic disk, with a break scale at $\ell_{0}^{-1}$, with $k^{-1/2}$ on low $k$ and Grete et al. (2021)’s $k^{-4/3}$ at high $k$.

### v2

@@ 4.1 Definitions
$\frac{df^{2}(\bm{k})}{d\bm{k}}=\tilde{\bm{f}}(\bm{k})\cdot\tilde{\bm{f}}^{\dagger}(\bm{k}),$ (22)
where $\square^{\dagger}$ denotes the complex conjugate. Even though there is a global anisotropy imparted on the flow due to $\bm{\nabla}\phi$, it is never strong enough to make the $\ell\parallel\bm{\nabla}\phi$ and $\ell\perp\bm{\nabla}\phi$ statistics significantly different (Beattie et al., 2025c). Hence, in this study we integrate the three-dimensional spectrum isotropically over solid-angle shells $\operatorname{d}\!{\Omega_{k}}$, where $k=|\bm{k}|$,
$\frac{df^{2}(k)}{dk}=\int\operatorname{d}\!{\Omega_{k}}\;k^{2}\frac{df^{2}(\bm{k})}{d\bm{k}},$ (23)
resulting in the one-dimensional isotropic spectrum, $df^{2}(k)/dk$.
@@ 4.2 The $k$-space structure of the turbulence
In Figure 7, we plot the integrated spectra of the total velocity, $du^{2}(k)/dk$ (top row), the kinetic energy $d\rho u^{2}(k)/dk$ (second row), the decomposed compressible $du_{c}^{2}(k)/dk$ and incompressible $du_{s}^{2}(k)/dk$ velocity modes (third row), and the sound speed $dc_{s}^{2}(k)/dk$ (bottom row), for MW (left column) and SB (right column), all normalized by the integral power. The gaseous scale height, $k_{0}\sim\ell_{0}^{-1}$, is indicated with the grey vertical line. All fitted indexes for spectra are shown in Table 3, and the exact empirical indexes, with the caveats associated with measuring over a fixed range of $k$ in simulations that are not scale-free, are discussed in Appendix B.
@@
Beginning with the first row, we compensate the velocity spectra for Kolmogorov (1941)-type turbulence, and we find that all models have a shallower than the Kolmogorov (1941) spectrum, and MW models follow the $k^{-3/2}$ spectrum found in the magnetohydrodynamic (MHD) SNe-driven turbulence study of Gent et al. (2021) and also the MHD studies by Beattie et al. (2025a) and Iroshnikov (1964). This spectrum has also been derived for weak acoustic wave turbulence (Zakharov & Sagdeev, 1970; Kochurin & Kuznetsov, 2024), and shown to potentially be embedded within supersonic turbulence if one completely resolves a subsonic cascade (Ferrand et al., 2020; Beattie et al., 2025a). This is plausible for our set up if SNe shocks become linear sound waves on small enough scales.
@@
Once we decompose each simulation into $\bm{u}_{c}$ and $\bm{u}_{s}$ velocity modes (second row), we observe that, as found in a number of SNe-driven ISM simulations, e.g., Balsara et al. (2004), Padoan et al. (2016), Pan et al. (2016) and Beattie et al. (2025c), the turbulence is strongly dominated by the incompressible $\bm{u}_{s}$ modes, with a spectra shallower than $k^{-5/3}$, and especially so for the high-SN rate SB model. This aligns also with work that found incompressible mode dominance to be true even for purely compressible-driven simulations (Federrath, 2013), but likely the mechanism for generating the incompressible modes varies significantly between the simulations, as discussed in Beattie et al. (2025c), i.e., the shallower spectrum dominates at smaller scales than the scale height, even more so than in purely incompressible-driven turbulent boxes (Federrath, 2013). Beattie et al. (2025c) proposed that phase mixing through the fractal cooling layer of SNRs is the strongest source of incompressible turbulence in SNe-driven turbulence.
@@
The compressible mode spectrum in the second row closely follows Burgers (1948) $k^{-2}$ prediction for strongly nonlinear compressible turbulence for the models, with some shallowing towards $k^{-5/3}$ in the SBFX model, but generally, $k^{-2}$ without a break scale. We also plot the pseudo-acoustic wave spectrum $k^{-3}$, which has been realized in low-$\mathcal{M}$ ($\mathcal{M}\lesssim 0.2$) hydrodynamical turbulence (Wang et al., 2017; Beattie et al., 2025b), but as we see here, none of the simulations are in this regime.
@@
Because the turbulence is dominated by $\bm{u}_{s}$ modes it is unlikely that the $k^{-3/2}$ $\bm{u}$ spectrum comes from the weak acoustic regime, and in order to get shallower spectra analogous to MHD turbulence (like $k^{-3/2}$) there must be a reduction in the turbulent nonlinearity (Boldyrev, 2006; Matthaeus et al., 2008; Chernoglazov et al., 2021; Beattie & Bhattacharjee, 2025), which in our case can possibly be attributed to the strong alignment between velocity $\bm{u}$ and vorticity $\bm{\omega}$ (the kinetic helicity), as seen in Beattie et al. (2025c) (in their Appendix, Figure 21) and previously in (Käpylä et al., 2018). The presence of an inverse cascade can also cause a shallower spectrum, as shown in detail by Beattie et al. (2025c) which is shown to be possible in helical, hydrodynamical turbulence (Plunian et al., 2020), potentially the kind of turbulence that is generated on the expanding SNRs.
@@
In the third row we plot the kinetic energy spectra compensated by $k^{-4/3}$, found previously in compressible (magnetized) turbulence (Grete et al., 2021). For $k\lesssim\ell_{0}^{-1}$, we see close alignment between the $k^{-4/3}$ model and $d\rho u^{2}(k)/dk$. For SB, we see this behavior at small and large scales, but with a transition between at $k\approx\ell_{0}^{-1}$. This transition is most likely due to the gaseous profile of the disk imprinting itself across the spectrum. As discussed in the Appendix of Beattie et al. (2025c), this increase in kinetic energy density at high $k$, compared to $du^{2}(k)/dk$, is likely due to high-density clumps on small scales. The velocity power spectrum is more universal across scales and between models, suggesting the velocity modes are correlated through the turbulence in the winds and disk, which we will explore in more detail in Section 4.4.
@@ 4.3 What sets the large-scale phase structure in the ISM?
which means that, in the adiabatic limit where $t_{\rm turb}\ll t_{\rm cool}$, $dc_{s}^{2}(k)/dk\sim du_{c}^{2}(k)/dk$, as we show in the bottom row [^2] In Appendix C we also show the relation between $du_{c}^{2}(k)/dk$ and $d(\bm{\nabla}\cdot\bm{u})^{2}(k)/dk$, i.e., $du_{c}^{2}(k)/dk\sim d(\bm{\nabla}\cdot\bm{u})^{2}(k)/(k^{2}dk)\sim dc_{s}^{2}(k)/dk\propto k^{-2}$. of Figure 7. It should now make sense that $dc_{s}^{2}(k)/dk$ for the MW model follows almost exactly the $\bm{u}_{c}$ spectrum, given that the disk has a strong adiabatic component (Figure 6). Any slight deviations are likely due to the existence of non-adiabatic thermodynamic processes. For the wind-dominated SB model, at the disk scale and onward ($\ell_{0}$ annotated in black dashed line), the spectrum largely follows $k^{-2}$, since the SB disk is also largely adiabatic. For $k\lesssim\ell^{-1}_{0}$ the spectrum deviates, because the winds are not principally adiabatic and have a mix of thermodynamic processes that we do not account for in our simple adiabatic model. This shows the the important role of compressible turbulent modes in controlling the temperature and phases of the ISMs in our models.
Figure 8: One-dimensional, isotropic velocity (top) and kinetic energy (bottom) power spectrum, normalized into global gaseous scale heights, $k\ell_{0}/2\pi=1$, and power on the scale height (e.g., $du^{2}(k)/du^{2}(k\ell_{0}/2\pi)=1$) for each galaxy model (MW; dark purple, SB; light purple). We show different SNe seeding schemes (SC; dashed linestyle, FX; solid linestyle). The data that is greyed is either MW, where $k<\ell_{0}^{-1}$, or SB $k>\ell_{0}^{-1}$. By normalizing the spectrum into the frame of the scale height we find there is a smooth transition between both the velocity and kinetic energy for the turbulence in the winds and the turbulence in the disk. The velocity spectrum exhibits a universal power law, $du^{2}(k)/dk\propto k^{-3/2}$, regardless of galaxy model or seeding prescription, whereas the kinetic energy spectrum retains some structure of the galactic disk, varying from $k^{-1/2}$ on $k<\ell_{0}^{-1}$ to $k^{-4/3}$ on $k>\ell_{0}^{-1}$. We indicate the global outer scale $\ell_{\rm{cor}}$ of the $du^{2}(k)/dk$ and $d\rho u^{2}(k)/dk$ spectra (Equation 28) in the purple shaded line, showing (particularly for the $du^{2}(k)/dk$ spectrum) that even though we inject momentum from SNe on the smallest scales in the simulations ($\ell_{\rm inj}=2\rm{dx}$), they drive turbulence to scales well beyond $\ell_{0}$, with $\ell_{\rm{cor}}\approx 6\ell_{0}$.
@@ 4.4 Cascading from the winds to the disk
Based on Figure 7, it is clear that many of the spectra seem to follow similar power law scalings across the different seeding and galactic models (which differ in mass, potential depth and injection energy). To explore this universality further, we transform the velocity and kinetic energy spectra such that $k\ell_{0}/2\pi=1$ for all models, and $du^{2}(k)/du^{2}(k\ell_{0}/2\pi)=1$, such that all models are placed in a global gaseous scale height reference frame. We plot the transformed spectra in Figure 8, showing the SB simulations in light purple (mostly situated in $k\lesssim\ell^{-1}_{0}$ modes, corresponding to the wind scales), and MW simulations in dark purple (mostly on the $k\gtrsim\ell^{-1}_{0}$ scales), and the two different seeding schemes with the different linestyles. We change the color of the spectra for the MW models on $k<\ell_{0}^{-1}$ using grey lines to indicate the spectra beyond those modes, and similarly for the SB spectra but for $k>\ell_{0}^{-1}$. This allows us to directly see how continuous the spectra are across $k\ell_{0}/2\pi$, whilst highlighting the structure of the turbulence in the different simulations below and above $k\ell_{0}/2\pi$.
We find that all spectra collapse across galactic models and seeding schemes into just a single velocity (top) and kinetic (bottom) spectrum, that are smooth across $\ell^{-1}_{0}$. The $du^{2}(k)/dk\propto k^{-3/2}$ discussed in Section 4.2 and previously measured in Padoan et al. (2016) and Beattie et al. (2025c) extends across both simulations, connecting the winds to the turbulent disk in a continuous manner. For the $d\rho u^{2}(k)/dk$ spectrum we find a broken power law, that seems to break at roughly $k\ell_{0}/2\pi\approx 1$, with $d\rho u^{2}(k)/dk\propto k^{-1/2}$ on $k\ell_{0}/2\pi\lesssim 1$ and $d\rho u^{2}(k)/dk\propto k^{-4/3}$ on $k\ell_{0}/2\pi\gtrsim 1$. This is likely due to the gaseous disk imprinting itself on the power spectrum, due to the mass density contribution in the spectrum. Federrath (2013) previously found that the $d\rho u^{2}(k)/dk$ was sensitive to the underlying driving field and $\mathcal{M}$, whereas the $du^{2}(k)/dk$ was not.
@@
$\frac{\ell_{\mathrm{cor}}}{\ell_{0}}=\frac{1}{\left\langle f^{2}\right\rangle}\displaystyle\int_{0}^{\infty}dk\;(k\ell_{0}/2\pi)^{-1}\frac{df^{2}(k)}{dk},$ (28)
in the global $\ell_{0}^{-1}$ frame using definitions from Beattie et al. (2025c), and find $\ell_{\rm cor}\approx 6\ell_{0}$ for $du^{2}(k)/dk$, showing that the largest scale velocity correlations extend well beyond the gaseous scale-height, and $\ell_{\rm cor}\approx 3\ell_{0}$ for $d\rho u^{2}(k)/dk$, being truncated at small scales due to the imprint of the galactic disk on $d\rho u^{2}(k)/dk$. Both the spectra (but particularly the velocity spectrum), show that SNe can drive turbulence to large scales, larger than the conventional wisdom, $\ell_{0}\approx\ell_{\rm cor}$ (Haverkorn et al., 2008; Liu et al., 2022; Beattie et al., 2022a).
@@ 5 Summary and Conclusions
  • For each of the models we calculate the power spectrum of the total velocity, kinetic energy, Helmholtz decomposed velocity modes, and sound speed and plot them Figure 7. In the disk-dominated MW simulations we show the entire phase structure of the predominantly adiabatic disk can be explained by the $\bm{u}_{c}$ mode turbulence ($dc_{s}^{2}(k)/dk\propto du_{c}^{2}(k)/dk\propto k^{-2}$), which we can derive by relating $c_{s}$ and $u_{c}$ in the weak-cooling, adiabatic limit, assuming that the compressible modes are sourced from turbulence within sound waves, detailed in Equations 24-27. Consistent with Beattie et al. (2025c), we find a velocity spectrum that is dominated by incompressible modes and close to $du^{2}(k)/dk\propto k^{-3/2}$, suggesting a weakening of the turbulent nonlinearity which is robust to the different simulation parameters.
  • By transforming both the velocity $u$ and kinetic energy $\rho u^{2}$ spectrum into units of $\ell_{0}$ and power on $\ell_{0}$, in Figure 8 we show that the wind-dominated (starburst; SB) and disk-dominated (Milky Way; MW) galaxy models have power spectra that smoothly transition into one another, suggesting the structure of the turbulence is robust to changes in the galactic potential, SNe-driving rate, and seeding prescription. Furthermore, a universal $du^{2}(k)/dk\propto k^{-3/2}$ spectrum extends between the disks and the winds seamlessly, suggesting a wind-disk turbulence connection, with the velocity cascade moving through the different ISM phases without hindrance. Furthermore, in this frame, we show that the outer scale of the velocity is $\approx 6\ell_{0}$, challenging the conventional wisdom that the outer scale is truncated at the gaseous scale height of the disk. The kinetic energy spectrum traces large-scale structure from the galactic disk, with a break scale at $\ell_{0}^{-1}$, with $k^{-1/2}$ on low $k$ and approximately Grete et al. (2021)’s $k^{-4/3}$ at high $k$. Higher resolution simulations will be required for definitively determining the spectral indexes, but we stress there is no evidence for a Kolmogorov-type $k^{-5/3}$ scaling in either the velocity of kinetic energy spectra.